### v1

@@ 4 Continuity of the supremum function $\Sigma C(X)\times\Sigma C(X)\rightarrow\Sigma C(X)$
$s_{n}(x_{1},x_{2},\dots,x_{n})=\ \downarrow\!{\{x_{1},x_{2},\dots,x_{n}\}}.$
@@
[14, Proposition 1.1] A space is an essentially complete $T_{0}$ space iff the following are satisfied:
@@
    The associated pre-order $(X,\leq)$ is a complete lattice.
@@
Hoffmann noticed that the above condition $(iii)$ can be replaced by that “$\sup\colon X\times X\rightarrow X$ is continuous” [14, Lemma 1.5].
Proposition 4.19.
Let $X,Y,Z$ be monotone determined spaces. A map $f:X\otimes Y\to Z$ is continuous iff it is separately continuous.
@@
$(2)\implies(1)$. For the completeness, we give the detailed proof here. Assume that $s_{n}$ is continuous for all $n\in\mathbb{N}$. Let $\mathcal{U}$ be an open subset of $\Sigma(C(X))$ and $A\in\mathcal{U}$. Without loss of generality, we assume $A\not=\emptyset$. Note that since $A=\bigcup\{\downarrow\!F:F\subseteq_{f}\!A\}$ and $\{\downarrow F:F\subseteq_{f}A\}$ is a directed family in $C(X)$, there exists a non-empty finite subset $F$ of $A$ such that $\downarrow\!F\in\mathcal{U}$. Let $F=\{x_{1},x_{2},\ldots,x_{n}\}$, then $s_{n}(x_{1},x_{2},\ldots,x_{n})=\downarrow\!F\in\mathcal{U}$. It follows that $(x_{1},x_{2},\ldots,x_{n})\in s_{n}^{-1}(\mathcal{U})$. By the continuity of $s_{n}$, there exists a family of open subsets $U_{k}(1\leq k\leq n)$ such that $U_{1}\times U_{2}\times\cdots\times U_{n}$ is open in $\prod\limits^{n}X$ and
@@
Since $x_{k}\in A$ for $1\leq k\leq n$, we have $A\in\Diamond U_{k}=\{B\in C(X):\ B\cap U_{k}\not=\emptyset\}$. It follows that $A\in\bigcap\limits_{k=1}^{n}\Diamond U_{k}\in\upsilon(C(X))$. For any $B\in\bigcap\limits_{k=1}^{n}\Diamond U_{k}$, there exists $y_{k}\in B\cap U_{k}$ for $1\leq k\leq n$. Since $(y_{1},y_{2},\dots,y_{n})\in s_{n}^{-1}(\mathcal{U})$, we have $\bigcup\limits_{k=1}^{n}\downarrow y_{k}\in\mathcal{U}$. It follows that $B\in\mathcal{U}$, i.e., $A\in\bigcap\limits_{k=1}^{n}\Diamond U_{k}\subseteq\mathcal{U}$. ∎
@@
Let $X$ be a monotone determined space. We consider a topology $\tau$ on $C(X)$ with $v(C(X))\subseteq\tau\subseteq\sigma(C(X))$. By a similar argument in Theorem 4.20, we can show that $\tau$ is equal to $v(C(X))$ if the operator $\bigcup\colon(C(X),\tau)\times(C(X),\tau)\rightarrow(C(X),\tau)$ is continuous. In other words, there exists only one topology $\tau$ on $C(X)$ such that $(1)$ the specialization order of $\tau$ is equal to the inclusion order and $(2)$ $(C(X),\tau)$ is an essentially complete $T_{0}$ space. But this is false for a non-monotone determined space $X$, for instance see Example 4.22.
@@
[4, Proposition 3.5] Let $P$ be a poset. If ${\displaystyle\Sigma(\prod^{n}P)=\prod^{n}(\Sigma P)}$ for each $n\in\mathbb{N}$, then $\sigma(\Gamma(P))=v(\Gamma(P))$.
@@
$(2)$ implies $(1)$ can be found in [4, Proposition 3.15].The above results does not hold for $T_{0}$ spaces. For example we consider $X=\Sigma L$ where $L$ is the Isbell’s lattice, then $(3)$ holds while $(2)$ does not hold since $\Sigma C(X)$ is not sober [21].
@@
There exists a non-monotone determined space $X$ such that $\eta\colon X\rightarrow\Sigma C(X)$ is continuous and $\sigma(C(X))=v(C(X))$. We consider the space $X$ in Example 3.17, it follows that $C(X)\cong\Gamma(D)$ since $\Sigma D$ is the D-completion of $X$. Notice that ${\displaystyle Id(\prod^{n}D)=\prod^{n}Id(D)}$, by using Corollary 4.26 and Proposition 4.29 we obtain that $\sigma(\Gamma(D))=v(\Gamma(D))$. Hence we also have $\sigma(C(X))=v(C(X))$. It is easy to see that $\eta\colon X\rightarrow\Sigma(C(X))$ is continuous.
@@
By the above lemma, if $\Sigma L\times\Sigma M$ is not a Scott space, then ${}\bigvee\colon\Sigma S\times\Sigma S\rightarrow\Sigma S$ is not continuous. We recall that a topological space $X$ is called a Scott space if its topology is equal to the Scott topology on $(X,\leq)$, which $\leq$ is the specialization order of $X$.
@@
[11, Proposition II-3.15(iii)] If $X$ is a continuous retract of a monotone convergence space $Y$, then $X$ must be $T_{0}$ and is a monotone convergence space. In the special case that $Y$ is a dcpo equipped with the Scott topology, then $X$ must also have the Scott topology.
@@
We use Hertling’s example [13, Theorem 7.4] to show that there exists a complete lattice $Z$ such that condition $(2)$ holds but condition $(1)$ fails. $Z$ is the disjoint union of a complete lattice $X$ and $Y=\sigma(X)$ plus a top element and a bottom element, i.e., $Z=\{X+Y\}\bigcup\{\top,\bot\}$. $Z$ has the following conditions:
@@ 5 Quasicontinuity of the lattice of closed subsets
We introduce the concepts of quasicontinuous spaces and quasialgebraic spaces. Then we will summarize the correspondences among the continuity of a $T_{0}$ space, the continuity of the lattice of its open subsets, and the continuity of the lattice of its closed subsets.
Given any set $X$, we use $F\subseteq_{f}X$ to denote that $F$ is a finite subset of $X$. Given any two subsets $G,H\subseteq X$, we define $G\leq H$ iff $\uparrow H\subseteq\uparrow G$. A family of finite sets is said to be directed if given $F_{1},F_{2}$ in the family, there exists an $F$ in the family such that $F_{1},F_{2}\leq F$. Let $X$ be a $T_{0}$ topological space and $\mathcal{M}$ be a directed family of finite subsets of $X$. Then $F_{\mathcal{M}}=\{A\subseteq X:\exists M\in\mathcal{M},\uparrow M\subseteq A\}$ is a filter under the inclusion order. We say that $\mathcal{M}$ converges to $x$ if $F_{\mathcal{M}}$ converges to $x$, i.e., for any open neighbourhood $U$ of $x$, there exists some $A\in\mathcal{M}$ such that $A\subseteq U$.
@@
Let $X$ be a $T_{0}$ space and $G,H\subseteq X$. We say that $G$ $n$-$approximates$ $H$, denoted by $G\ll_{n}H$, if for every net $(x_{j})_{J}$ of $X$, $(x_{j})_{J}\to y$ for some $y\in H$ implies $(x_{j})_{J}\cap\uparrow G\neq\emptyset$. We write $G\ll_{n}x$ for $G\ll_{n}\{x\}$. $G$ is said to be $n$-compact if $G$ is a finite set and $G\ll_{n}G$.
@@
For any $T_{0}$ space $X$, the following are equivalent: for all $x,y\in Y$,
@@
    $H\subseteq(\uparrow G)^{\circ}$, where $(\uparrow G)^{\circ}$ means the interior of $\uparrow G$.
Proof. $(2)\Rightarrow(1)$ is obvious. We only show $(1)\Rightarrow(2)$. Assume that $y\in H,y\not\in(\uparrow G)^{\circ}$. Then, for any open set $U\subseteq X$ with $y\in U$, there exists a $y_{U}\in U$ such that $y_{U}\not\in\uparrow G$. Set $J=\{U\in\mathcal{O}(X):y\in U\}$ ordered as: $\forall U,V\in J$,$U\leq V\Leftrightarrow V\subseteq U.$ Then $J$ is directed. Easily, one can see that $(y_{U})_{U\in J}$ is a net converging to $y$. Since $G\ll_{n}y$, there exists $U_{0}\in J$ such that $x\sqsubseteq y_{U_{0}}$ for some $x\in G$ by definition. It is a contradiction for $y_{U_{0}}\not\in\uparrow G$. Hence, $y\in(\uparrow G)^{\circ}$.  $\Box$
Since the $n$-approximation does not depend on monotone determined space, a notion of $n$-quasicontinuity for a $T_{0}$ space can be defined as follows.
@@
A $T_{0}$ topological space $X$ is called $n$-quasicontinuous if for each $x\in X$, there is a directed family $D(x)\subseteq fin_{n}(x)=\{F:F\ \text{is finite},F\ll_{n}x\}$ such that $D(x)\rightarrow x$.
@@
A $T_{0}$ topological space $X$ is called locally hypercompact (or finitary compact), if for any open subsets $U$ and $x\in U$, there exists some $F\subseteq_{f}X$ such that $x\in(\uparrow F)^{\circ}\subseteq\uparrow F\subseteq U$. $X$ is called hypercompactly based if for any open subsets $U$ and $x\in U$, there exists some $F\subseteq_{f}X$ such that $x\in(\uparrow F)^{\circ}=\uparrow F\subseteq U$.
Since $G\ll_{n}H$ iff $H\subseteq(\uparrow G)^{\circ}$, definitions of $n$-quasicontinuous and locally hypercompact are equivalent. Similarly, we can define $n$-quasialgebraic spaces. And the following statements hold.
@@
[9, 19] For a $T_{0}$ topological space $X$, the following are equivalent:
@@
A $T_{0}$ topological space is called quasicontinuous if one of the four equivalent conditions in Theorem 5.38 holds; called quasialgebraic if it is $n$-quasialgebraic.
@@
[4, Lemma 2.3] Let $X$ be a topological space and $U,V\in\mathcal{O}(X)$. $U\prec V$ iff $U\subseteq\uparrow F\subseteq V$ for some $F\subseteq_{f}X$.
@@
Lemma 5.41 does not always hold for general $T_{0}$ spaces. A simple example is $\mathbb{N}$, the set of natural numbers, endowed with the cofinite topology. Any open subset of it is compact, but not hypercompact.
By the above discussion, we conclude the relationships between the continuity of a $T_{0}$ space and the lattice of its open subsets as follows.
@@
Given any $T_{0}$ space $X$, denote $C(X)$ the lattice of its closed subsets. We show that the quasicontinuity of $X$ can be characterized by the quasicontinuity of $C(X)$, and vice versa. We first recall some equivalent conditions for continuity of lattices.
@@
Let $X$ be a $T_{0}$ space. If $C(X)$ is a quasicontinuous lattice, then $X$ is quasicontinuous.
@@
Let $C(X)$ be a quasicontinuous domain. Then for any $x\in X$, there exists some $\{A_{1},\dots,A_{n}\}\ll\ \downarrow x$ in $C(X)$. Given any $a_{i}\in A_{i}\ (1\leq i\leq n)$, we have $\{\downarrow a_{1},\dots,\downarrow a_{n}\}\ll\ \downarrow x$.
@@
Given any net $(x_{i})_{I}\to x$, $x\in\overline{\bigcup_{i\in I}\{x_{i}\}}$. Then $\downarrow x\subseteq\ \overline{\bigcup_{i\in I}\{x_{i}\}}$. Let =$\mathcal{F}=\{\bigcup_{i\in F}\!\downarrow x_{i}:F\subseteq_{f}I\}$. Then $\mathcal{F}$ is directed and $\downarrow x\leq\bigvee^{\uparrow}\mathcal{F}=\overline{\bigcup_{i\in I}\{x_{i}\}}$ in $C(X)$. There exist some $F\subseteq_{f}I$ and $a_{j}$ such that $\downarrow a_{j}\subseteq\bigcup_{i\in F}\downarrow x_{i}$. Thus, there exists a $i\in F$ such that $a_{j}\leq x_{i}$, i.e., $\{a_{1},\dots,a_{n}\}\ll_{n}x$ in $X$.
Claim 2. For any $x$ and any open neighbourhood $U$ of $x$, there exists some $F\subseteq_{f}U$ such that $x\in(\uparrow F)^{\circ}$.
Given any directed family $\mathcal{A}=\{A_{i}\}_{I}$ of closed subsets of $X$ such that $\downarrow\!x\leq\bigvee\mathcal{A}=\overline{\bigcup_{i\in I}A_{i}}$, then $U\cap(\bigcup_{i\in I}A_{i})\not=\emptyset$. There exists some $i\in I$ such that $A_{i}\cap U\not=\emptyset$. Pick $b\in A_{i}\cap U$, then $\downarrow\!b\subseteq A_{i}$. We have $\{\downarrow\!b:b\in U\}\ll\ \downarrow\!x$ in $C(X)$. Since $C(X)$ is quasicontinuous, there exists a $\{B_{1},\dots,B_{m}\}\subseteq C(X)$ such that $\{\downarrow b:b\in U\}\ll\{B_{1},\dots,B_{m}\}\ll\ \downarrow x$ in $C(X)$. Then
$\forall 1\leq i\leq m,\ \{\downarrow b:b\in U\}\ll B_{i}=\overline{\bigcup_{F\subseteq_{f}B_{i}}\downarrow F},$
which implies that there exists some $F\subseteq_{f}B_{i}$ and $b_{i}\in U$ such that ${\downarrow}b_{i}\subseteq{\downarrow}F\subseteq B_{i}$. Thus, $b_{i}\in B_{i}\cap U$. It follows that $\{{\downarrow}b_{1},\dots,{\downarrow}b_{m}\}\ll\ {\downarrow}x$ in $C(X)$. Then $\{b_{1},\dots,b_{n}\}\ll_{n}x$ by Claim 1, i.e., $x\in({\uparrow}\{b_{1},\dots,b_{n}\})^{\circ}$. ∎
@@
Let $X$ be a $T_{0}$ topological space. Then the following conditions are equivalent:
  • (1)
@@
  • (2)
    $\mathcal{O}(X)$ with the upper topology is an injective space;
  • (3)
@@
  • (4)
    $C(X)$ with Lawson topology is Hausdorff;
  • (5)

### v2

@@ 4 Continuity of the supremum function $\Sigma C(X)\times\Sigma C(X)\rightarrow\Sigma C(X)$
$s_{n}(x_{1},x_{2},\dots,x_{n})=\ \downarrow\hskip-2.0pt\!{\{x_{1},x_{2},\dots,x_{n}\}}.$
@@
[14, Proposition 1.1] A space is an essentially complete $\mathbf{T_{0}}$ space iff the following are satisfied:
@@
    The associated pre-order (i.e., specialization order) $(X,\leq)$ is a complete lattice.
@@
Hoffmann noticed that the above condition $(iii)$ can be replaced by that $\sup\colon X\times X\rightarrow X$ is continuous” [14, Lemma 1.5].
@@
$(2)\implies(1)$. For the completeness, we give the detailed proof here. Assume that $s_{n}$ is continuous for all $n\in\mathbb{N}$. Let $\mathcal{U}$ be an open subset of $\Sigma(C(X))$ and $A\in\mathcal{U}$. Without loss of generality, we assume $A\not$$\emptyset$. Note that since $A=\bigcup\{\downarrow\hskip-2.0pt\!F:F\subseteq_{f}\!A\}$ and $\{\downarrow\hskip-2.0ptF:F\subseteq_{f}A\}$ is a directed family in $C(X)$, there exists a non-empty finite subset $F$ of $A$ such that $\downarrow\hskip-2.0pt\!F\in\mathcal{U}$. Let $F=\{x_{1},x_{2},\ldots,x_{n}\}$, then $s_{n}(x_{1},x_{2},\ldots,x_{n})=\downarrow\hskip-2.0pt\!F\in\mathcal{U}$. It follows that $(x_{1},x_{2},\ldots,x_{n})\in s_{n}^{-1}(\mathcal{U})$. By the continuity of $s_{n}$, there exists a family of open subsets $U_{k}(1\leq k\leq n)$ such that $U_{1}\times U_{2}\times\cdots\times U_{n}$ is open in $\prod\limits^{n}X$ and
@@
Since $x_{k}\in A$ for $1\leq k\leq n$, we have $A\in\Diamond U_{k}=\{B\in C(X):\ B\cap U_{k}\not$$\emptyset\}$. It follows that $A\in\bigcap\limits_{k=1}^{n}\Diamond U_{k}\in\upsilon(C(X))$. For any $B\in\bigcap\limits_{k=1}^{n}\Diamond U_{k}$, there exists $y_{k}\in B\cap U_{k}$ for $1\leq k\leq n$. Since $(y_{1},y_{2},\dots,y_{n})\in s_{n}^{-1}(\mathcal{U})$, we have $\bigcup\limits_{k=1}^{n}\downarrow\hskip-2.0pty_{k}\in\mathcal{U}$. It follows that $B\in\mathcal{U}$, i.e., $A\in\bigcap\limits_{k=1}^{n}\Diamond U_{k}\subseteq\mathcal{U}$. ∎
@@
Let $X$ be a monotone determined space. We consider a topology $\tau$ on $C(X)$ with $v(C(X))\subseteq\tau\subseteq\sigma(C(X))$. By a similar argument in Theorem 4.26, we can show that $\tau$ is equal to $v(C(X))$ if the operator $\bigcup\colon(C(X),\tau)\times(C(X),\tau)\rightarrow(C(X),\tau)$ is continuous. In other words, there exists only one topology $\tau$ on $C(X)$ such that $(1)$ the specialization order of $\tau$ is equal to the inclusion order and $(2)$ $(C(X),\tau)$ is an essentially complete $\mathbf{T_{0}}$ space. But this is false for a non-monotone determined space $X$, for instance see Example 4.28.
@@
[4, Proposition 3.5] Let $P$ be a poset. If $\Sigma(\prod^{n}P)=\prod^{n}(\Sigma P)$} for each $n\in\mathbb{N}$, then $\sigma(\Gamma(P))=v(\Gamma(P))$.
@@
$(2)$ implies $(1)$ can be found in [4, Proposition 3.15].The above results do not hold for $\mathbf{T_{0}}$ spaces. For example we consider $X=\Sigma L$ where $L$ is the Isbell’s lattice, then $(3)$ holds while $(2)$ does not hold since $\Sigma C(X)$ is not sober [21].
@@
There exists a non-monotone determined space $X$ such that $\eta\colon X\rightarrow\Sigma C(X)$ is continuous and $\sigma(C(X))=v(C(X))$. We consider the space $X$ in Example 3.16, it follows that $C(X)\cong\Gamma(D)$ since $\Sigma D$ is the D-completion of $X$. Notice that $Id(\prod^{n}D)=\prod^{n}Id(D)$}, by using Corollary 4.32 and Proposition 4.35 we obtain that $\sigma(\Gamma(D))=v(\Gamma(D))$. Hence we also have $\sigma(C(X))=v(C(X))$. It is easy to see that $\eta\colon X\rightarrow\Sigma(C(X))$ is continuous.
@@
By the above lemma, if $\Sigma L\times\Sigma M$ is not a Scott space, then $\bigvee\colon\Sigma S\times\Sigma S\rightarrow\Sigma S$ is not continuous. We recall that a topological space $X$ is called a Scott space if its topology is equal to the Scott topology on $(X,\leq)$, where $\leq$ is the specialization order of $X$.
@@
[11, Proposition II-3.15(iii)] If $X$ is a continuous retract of a monotone convergence space $Y$, then $X$ must be $\mathbf{T_{0}}$ and is a monotone convergence space. In the special case that $Y$ is a dcpo equipped with the Scott topology, then $X$ must also have the Scott topology.
@@
We use Hertling’s example [13, Theorem 7.4] to show that there exists a complete lattice $Z$ such that condition $(2)$ holds but condition $(1)$ fails. $Z$ is the disjoint union of a complete lattice $X$ and $Y=\sigma(X)$ plus a top element and a bottom element, i.e., $\{X+Y\}\bigcup\{\top,\bot\}. ZZ$ has the following conditions:
@@ 5 Quasicontinuity of the lattice of closed subsets
We introduce the concepts of quasicontinuous spaces and quasialgebraic spaces. Then we will summarize the correspondences among the continuity of a $\mathbf{T_{0}}$ space, the continuity of the lattice of its open subsets, and the continuity of the lattice of its closed subsets.
Given any set $X$, we use $F\subseteq_{f}X$ to denote that $F$ is a finite subset of $X$. Given any two subsets $G,H\subseteq X$, we define $G\leq H$ iff $\uparrow\hskip-2.0ptH\subseteq\uparrow\hskip-2.0ptG$. A family of finite sets is said to be directed if given $F_{1},F_{2}$ in the family, there exists an $F$ in the family such that $F_{1},F_{2}\leq F$. Let $X$ be a $\mathbf{T_{0}}$ topological space and $\mathcal{M}$ be a directed family of finite subsets of $X$. Then $F_{\mathcal{M}}=\{A\subseteq X:\exists M\in\mathcal{M},\uparrow\hskip-2.0ptM\subseteq A\}$ is a filter under the inclusion order. We say that $\mathcal{M}$ converges to $x$ if $F_{\mathcal{M}}$ converges to $x$, i.e., for any open neighbourhood $U$ of $x$, there exists some $A\in\mathcal{M}$ such that $A\subseteq U$.
@@
Let $X$ be a $\mathbf{T_{0}}$ space and $G,H\subseteq X$. We say that $G$ $n$-approximatesapproximates $H$, denoted by $G\ll_{n}H$, if for every net $(x_{j})_{J}$ of $X$, $(x_{j})_{J}\to y$ for some $y\in H$ implies $(x_{j})_{J}\cap\uparrow\hskip-2.0ptG\neq\emptyset$. We write $G\ll_{n}x$ for $\ll_{n}\{x\}. GG$ is said to be $n$-compact if $G$ is a finite set and $G\ll_{n}G$.
@@
For any $\mathbf{T_{0}}$ space $X$, the following are equivalent: for all $x,y\in Y$,
@@
    $H\subseteq(\uparrow\hskip-2.0ptG)^{\circ}$, where $(\uparrow\hskip-2.0ptG)^{\circ}$ means the interior of $\uparrow\hskip-2.0ptG$.
Proof. $(2)\Rightarrow(1)$ is obvious. We only show $(1)\Rightarrow(2)$. Assume that $y\in H,y\not\in(\uparrow\hskip-2.0ptG)^{\circ}$. Then, for any open set $U\subseteq X$ with $y\in U$, there exists a $y_{U}\in U$ such that $y_{U}\not\in\uparrow\hskip-2.0ptG$. Set $J=\{U\in\mathcal{O}(X):y\in U\}$ ordered as: $\forall U,V\in J$$U\leq V\Leftrightarrow V\subseteq U.$ Then $J$ is directed. Easily, one can see that $(y_{U})_{U\in J}$ is a net converging to $y$. Since $G\ll_{n}y$, there exists $U_{0}\in J$ such that $x\sqsubseteq y_{U_{0}}$ for some $x\in G$ by definition. It is a contradiction for $y_{U_{0}}\not\in\uparrow\hskip-2.0ptG$. Hence, $y\in(\uparrow\hskip-2.0ptG)^{\circ}$.  $\Box$
Since the $n$-approximation does not depend on monotone determined space, a notion of $n$-quasicontinuity for a $\mathbf{T_{0}}$ space can be defined as follows.
@@
A $\mathbf{T_{0}}$ topological space $X$ is called $n$-quasicontinuous if for each $x\in X$, there is a directed family $D(x)\subseteq fin_{n}(x)=\{F:F\ \text{is finite},F\ll_{n}x\}$ such that $D(x)\rightarrow x$.
@@
A $\mathbf{T_{0}}$ topological space $X$ is called locally hypercompact (or finitary compact), if for any open subsets $U$ and $x\in U$, there exists some $F\subseteq_{f}X$ such that $x\in(\uparrow\hskip-2.0ptF)^{\circ}\subseteq\uparrow\hskip-2.0ptF\subseteq U$. $X$ is called hypercompactly based if for any open subsets $U$ and $x\in U$, there exists some $F\subseteq_{f}X$ such that $x\in(\uparrow\hskip-2.0ptF)^{\circ}=\uparrow\hskip-2.0ptF\subseteq U$.
Since $G\ll_{n}H$ iff $H\subseteq(\uparrow\hskip-2.0ptG)^{\circ}$, definitions of $n$-quasicontinuous and locally hypercompact are equivalent. Similarly, we can define $n$-quasialgebraic spaces. And the following statements hold.
@@
[9, 19] For a $\mathbf{T_{0}}$ topological space $X$, the following are equivalent:
@@
A $\mathbf{T_{0}}$ topological space is called quasicontinuous if one of the four equivalent conditions in Theorem 5.44 holds; called quasialgebraic if it is $n$-quasialgebraic.
@@
[4, Lemma 2.3] Let $X$ be a topological space and $\in\mathcal{O}(X). UVU\prec V$ iff $U\subseteq\uparrow\hskip-2.0ptF\subseteq V$ for some $F\subseteq_{f}X$.
@@
Lemma 5.47 does not always hold for general $\mathbf{T_{0}}$ spaces. A simple example is $\mathbb{N}$, the set of natural numbers, endowed with the cofinite topology. Any open subset of it is compact, but not hypercompact.
By the above discussion, we conclude the relationships between the continuity of a $\mathbf{T_{0}}$ space and the lattice of its open subsets as follows.
@@
Given any $\mathbf{T_{0}}$ space $X$, denote $C(X)$ the lattice of its closed subsets. We will study the relations between $X$ and $C(X)$. We first recall some useful conditions for continuity and distributiveness of lattices.
@@
Let $X$ be a $\mathbf{T_{0}}$ space. If $C(X)$ is a continuous lattice, then $X$ is continuous.
@@
$C(X)$ is a completely distributive lattice by Lemma 5.50 since it has enough co-primes. It follows that $\mathcal{O}(X)$ is completely distributive. Hence $X$ is continuous by Theorem 5.49(1). ∎
The above result can be generalized to quasicontinuity. Whether an analogous result holds for core-compactness, however, is still unknown.
Proposition 5.53.
Let $X$ be a $\mathbf{T_{0}}$ space. If $C(X)$ is a quasicontinuous lattice, then $X$ is quasicontinuous.
Proof.
Let $C(X)$ be a quasicontinuous domain. Then for any $x\in X$, there exists some $\{A_{1},\dots,A_{n}\}\ll\ \downarrow\hskip-2.0ptx$ in $C(X)$. Given any $a_{i}\in A_{i}\ (1\leq i\leq n)$, we have $\{\downarrow\hskip-2.0pta_{1},\dots,\downarrow\hskip-2.0pta_{n}\}\ll\ \downarrow\hskip-2.0ptx$.
@@
Given any net $(x_{i})_{I}\to x$, $x\in\overline{\bigcup_{i\in I}\{x_{i}\}}$. Then $\downarrow\hskip-2.0ptx\subseteq\ \overline{\bigcup_{i\in I}\{x_{i}\}}$. Let $\mathcal{F}=\{\bigcup_{i\in F}\!\downarrow\hskip-2.0ptx_{i}:F\subseteq_{f}I\}$. Then $\mathcal{F}$ is directed and $\downarrow\hskip-2.0ptx\leq\bigvee^{\uparrow}\hskip-2.0pt\mathcal{F}=\overline{\bigcup_{i\in I}\{x_{i}\}}$ in $C(X)$. There exist some $F\subseteq_{f}I$ and $a_{j}$ such that $\downarrow\hskip-2.0pta_{j}\subseteq\bigcup_{i\in F}\downarrow\hskip-2.0ptx_{i}$. Thus, there exists a $i\in F$ such that $a_{j}\leq x_{i}$, i.e., $\{a_{1},\dots,a_{n}\}\ll_{n}x$ in $X$.
Claim 2. For any $x$ and any open neighbourhood $U$ of $x$, there exists some $F\subseteq_{f}U$ such that $x\in(\uparrow\hskip-2.0ptF)^{\circ}$.
Given any directed family $\mathcal{A}=\{A_{i}\}_{I}$ of closed subsets of $X$ such that $\downarrow\hskip-2.0pt\!x\leq\bigvee\mathcal{A}=\overline{\bigcup_{i\in I}A_{i}}$, then $U\cap(\bigcup_{i\in I}A_{i})\not$$\emptyset$. There exists some $i\in I$ such that $A_{i}\cap U\not$$\emptyset$. Pick $b\in A_{i}\cap U$, then $\downarrow\hskip-2.0pt\!b\subseteq A_{i}$. We have $\{\downarrow\hskip-2.0pt\!b:b\in U\}\ll\ \downarrow\hskip-2.0pt\!x$ in $C(X)$. Since $C(X)$ is quasicontinuous, there exists a $\{B_{1},\dots,B_{m}\}\subseteq C(X)$ such that $\{\downarrow\hskip-2.0ptb:b\in U\}\ll\{B_{1},\dots,B_{m}\}\ll\ \downarrow\hskip-2.0ptx$ in $C(X)$. Then
$\forall 1\leq i\leq m,\ \{\downarrow\hskip-2.0ptb:b\in U\}\ll B_{i}=\overline{\bigcup_{F\subseteq_{f}B_{i}}\downarrow\hskip-2.0ptF},$
which implies that there exists some $F\subseteq_{f}B_{i}$ and $b_{i}\in U$ such that ${\downarrow\hskip-2.0pt}b_{i}\subseteq{\downarrow\hskip-2.0pt}F\subseteq B_{i}$. Thus, $b_{i}\in B_{i}\cap U$. It follows that $\{{\downarrow\hskip-2.0pt}b_{1},\dots,{\downarrow\hskip-2.0pt}b_{m}\}\ll\ {\downarrow\hskip-2.0pt}x$ in $C(X)$. Then $\{b_{1},\dots,b_{n}\}\ll_{n}x$ by Claim 1, i.e., $x\in({\uparrow\hskip-2.0pt}\{b_{1},\dots,b_{n}\})^{\circ}$. ∎
@@
Finally we collect some equivalent characterizations of quasicontinuous space as follows.
@@
Let $X$ be a $\mathbf{T_{0}}$ topological space. Then the following conditions are equivalent:
  1. (1)
@@
  2. (2)
    ([19, Theorem 2(2)]) $\mathcal{O}(X)$ with the upper topology is an injective space;
  3. (3)
@@
  4. (4)
    ([11, Theorem VII-3.10]) $C(X)$ with Lawson topology is Hausdorff;
  5. (5)